# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 6.2 Constraints on flavor-violating Higgs couplings
$\hfil\displaystyle d_{\rm Hg}\over[{\rm ecm}]$ $\supset{(1.2\pm 5.8)\cdot 10^{-24}\over(2\times 0.125^{2})}{\rm Im}(y^{d}_{12}y^{d}_{21})+{(-13.0\pm 8.1)\cdot 10^{-25}\over(2\times 0.125^{2})}{\rm Im}(y^{d}_{13}y^{d}_{31})$ (133)
$+{(-6.5\pm 4.6)\cdot 10^{-28}\over(2\times 0.125^{2})}{\rm Im}(y^{d}_{23}y^{d}_{32})+{(-2.5\pm 1.5)\cdot 10^{-25}\over(2\times 0.125^{2})}{\rm Im}(y^{u}_{12}y^{u}_{21})\,,$
@@
${\omega_{\textrm{HfF}}\over[\textrm{mrad/s}]}\supset{(-28.5\pm 1.0)\cdot 10^{5}\over(2\times 0.125^{2})}{\rm Im}(y^{e}_{12}y^{e}_{21})+{(-55.2\pm 2.4)\cdot 10^{6}\over(2\times 0.125^{2})}{\rm Im}(y^{e}_{13}y^{e}_{31}).$ (134)
Similarly, one can find expressions for the other EDM observables. Only a specific combination of the Higgs coupling can be probed with EDMs. To get an estimate of the upper bounds on these couplings one can compare the above expressions with the experimental limits, e.g. for $\left|d_{\rm Hg}\right|<6.2\times 10^{-30}$ $e$ cm and $\omega_{\textrm{HfF}}<0.17$ mrad/s.
Figure 5: EDM and flavor constraints on the CP and flavor violating Higgs couplings.
We perform also a combined fit of FV Higgs couplings to all EDM observables using the current and future upper limits given in Tab. 2. For this purpose, we assume only one combination of the couplings to be non-zero at a time. The results are presented in Tab. 11 (column 2).
In column 3, we indicate the origin of the EDMs in terms of low-energy WET operators at $\mu=2$GeV. At the EW scale the scalar and vector operators are generated by integrating out the Higgs as shown in the matching relations (127). Some of these give leading contribution to the EDMs through direct four-fermion matrix elements (see (3.4.1)- (60)) for the light flavors $f,f^{\prime}=u,d,s$. 1-loop QCD and QED effects through the operator mixing onto dipoles are also present for the light flavors. However, for the heavier generations, $c$ and $b$, the contribution to the EDMs arises solely due to operator mixing at 1-loop level (see (80)-(82)). Apart from that, as shown in column 3, the Weinberg operator also contributes for the case of heavier generations due to threshold corrections of the $b$ and $c$ dipoles onto it at the $m_{b}$ and $m_{c}$ scales (see (89)-(89)).
@@
Im($y^{d}_{12}y_{21}^{d}$) $(8\pm 6)\cdot 10^{-8}$ $(7\pm 4)\cdot 10^{-9}$ $\big{[}{C}_{d\gamma(G)}\big{]}_{11,22},\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{1221},\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{1122}$
Im($y_{13}^{d}y_{31}^{d}$) $(0\pm 4)\cdot 10^{-8}$ $(4\pm 3)\cdot 10^{-10}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{d\gamma(G)}\big{]}_{11}$
Im($y_{23}^{d}y_{32}^{d}$) $(-2\pm 8)\cdot 10^{-5}$ $(10\pm 7)\cdot 10^{-7}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{d\gamma(G)}\big{]}_{22}$
Im($y_{12}^{u}y_{21}^{u}$) $(2\pm 17)\cdot 10^{-8}$ $(17\pm 13)\cdot 10^{-10}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{u\gamma(G)}\big{]}_{11}$
Im($y_{12}^{e}y_{21}^{e}$) $(-4\pm 6)\cdot 10^{-10}$ $(-4\pm 6)\cdot 10^{-11}$ $\big{[}{C}_{e\gamma}\big{]}_{11}$
Im($y_{13}^{e}y_{31}^{e}$) $(-2\pm 3)\cdot 10^{-10}$ $(-2\pm 3)\cdot 10^{-12}$ $\big{[}{C}_{e\gamma}\big{]}_{11}$
Table 11: Current and future EDM constraints on the combinations of the generic flavor violating Higgs couplings (column 2). In the column 3, we also indicate the origin of EDM contributions at $\mu=2$ GeV.
@@ 6.3 EDM constraints on flavor-conserving Higgs couplings
For the flavor-conserving Yukawas, we get contributions from the tree-level matching onto four-fermion, given in Eq. (127), and from the two-loop Barr-Zee diagrams, in Eq. (6.1). Using these one can directly use the master formulae to obtain the EDMs in terms of the FC Yukawas, similar to the expressions in (133) and (134).
@@
here $f$ can be equal to $f^{\prime}$. Therefore, for $f=f^{\prime}$ and $i=j$, one can also probe individual couplings (more precisely a combination of its Re and Im parts: $\textrm{Re}(y^{f}_{ii})\textrm{Im}(y^{f}_{ii})$) in an uncorrelated manner to other couplings, under the assumption that only a single coupling is turned on. All in all, the tree-level and two-loop BZ diagram effects can test very different parts of the low-energy Higgs effective coupling parameter space. In Tab. 12 we show upper bounds at $1\sigma$ level on the various combinations of the generic FC Higgs couplings.
@@
Im$(y_{11}^{d}y_{11}^{d})$ $(19\pm 14)\times 10^{-9}$ $(29\pm 14)\times 10^{-10}$ $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{1111},\big{[}{C}_{d\gamma(G)}\big{]}_{11}$
Im$(y_{22}^{d}y_{22}^{d})$ $(-3\pm 12)\times 10^{-4}$ $(14\pm 11)\times 10^{-6}$ $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{2222},\big{[}{C}_{d\gamma(G)}\big{]}_{22}$
Im$(y_{33}^{d}y_{33}^{d})$ $\big{[}{C}_{\widetilde{G}}\big{]}$
Im$(y_{11}^{d}y_{22}^{d})$ $(15\pm 11)\times 10^{-9}$ $(24\pm 12)\times 10^{-10}$ $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{1122},\big{[}{C}_{d\gamma(G)}\big{]}_{11,22}$
Im$(y_{11}^{d}y_{33}^{d})$ $(17\pm 13)\times 10^{-9}$ $(21\pm 21)\times 10^{-10}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{d\gamma(G)}\big{]}_{11}$
Im$(y_{22}^{d}y_{33}^{d})$ $(-1\pm 5)\times 10^{-4}$ $(6\pm 4)\times 10^{-6}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{d\gamma(G)}\big{]}_{22}$
Im$(y_{11}^{u}y_{11}^{u})$ $(-15\pm 12)\times 10^{-9}$ $(5\pm 9)\times 10^{-10}$ $\big{[}{C}_{uu}^{S1(8),RR}\big{]}_{1111},\big{[}{C}_{u\gamma(G)}\big{]}_{11}$
Im$(y_{22}^{u}y_{22}^{u})$ $(0\pm 6)\times 10^{-2}$ $\big{[}{C}_{\widetilde{G}}\big{]}$
Im$(y_{11}^{u}y_{22}^{u})$ $(-15\pm 12)\times 10^{-9}$ $(4\pm 11)\times 10^{-10}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{u\gamma(G)}\big{]}_{11}$
Im$(y_{11}^{u}y_{11}^{d})$ $(-18\pm 17)\times 10^{-9}$ $(6\pm 5)\times 10^{-10}$ $\big{[}{C}_{ud}^{S1(8),RR}\big{]}_{1111},\big{[}{C}_{u\gamma(G)}\big{]}_{11},\big{[}{C}_{d\gamma(G)}\big{]}_{11}$
Im$(y_{11}^{u}y_{22}^{d})$ $(-10\pm 8)\times 10^{-9}$ $(3\pm 6)\times 10^{-10}$ $\big{[}{C}_{ud}^{S1(8),RR}\big{]}_{1122},\big{[}{C}_{u\gamma(G)}\big{]}_{11},\big{[}{C}_{d\gamma(G)}\big{]}_{22}$
Im$(y_{11}^{u}y_{33}^{d})$ $(-16\pm 12)\times 10^{-9}$ $(-9\pm 18)\times 10^{-10}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{u\gamma(G)}\big{]}_{11}$
Im$(y_{22}^{u}y_{11}^{d})$ $(17\pm 13)\times 10^{-9}$ $(4\pm 16)\times 10^{-10}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{d\gamma(G)}\big{]}_{11}$
Im$(y_{22}^{u}y_{22}^{d})$ $(-2\pm 8)\times 10^{-4}$ $(9\pm 7)\times 10^{-6}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{d\gamma(G)}\big{]}_{22}$
Im$(y_{22}^{u}y_{33}^{d})$ $\big{[}{C}_{\widetilde{G}}\big{]}$
Im$(y_{11}^{e}y_{11}^{e})$ $(-2\pm 3)\times 10^{-11}$ $(-2\pm 3)\times 10^{-12}$ $\big{[}{C}_{ee}^{S,RR}\big{]}_{1111},\big{[}{C}_{e\gamma}\big{]}_{11}$
Im$(y_{11}^{e}y_{22}^{e})$ $(-2\pm 3)\times 10^{-11}$ $(-2\pm 3)\times 10^{-12}$ $\big{[}{C}_{e\gamma}\big{]}_{11}$
Im$(y_{11}^{e}y_{33}^{e})$ $(-2\pm 3)\times 10^{-11}$ $(-2\pm 3)\times 10^{-12}$ $\big{[}{C}_{e\gamma}\big{]}_{11}$
Im$(y_{11}^{e}y_{11}^{d})$ $(-2\pm 3)\times 10^{-11}$ $(-2\pm 3)\times 10^{-12}$ $\big{[}{C}_{ed}^{S,RR}\big{]}_{1111},\big{[}{C}_{e(d)\gamma}\big{]}_{11}$
Im$(y_{11}^{e}y_{22}^{d})$ $(-2\pm 3)\times 10^{-11}$ $(-2\pm 3)\times 10^{-12}$ $\big{[}{C}_{ed}^{S,RR}\big{]}_{1122},\big{[}{C}_{e(d)\gamma}\big{]}_{11}$, $\big{[}{C}_{d\gamma(G)}\big{]}_{22}$
Im$(y_{11}^{e}y_{33}^{d})$ $(-2\pm 3)\times 10^{-11}$ $(-2\pm 3)\times 10^{-12}$ $\big{[}{C}_{e(d)\gamma}\big{]}_{11}$, $\big{[}{C}_{\tilde{G}}\big{]}$
Im$(y_{11}^{e}y_{11}^{u})$ $(-2\pm 3)\times 10^{-11}$ $(-2\pm 3)\times 10^{-12}$ $\big{[}{C}_{eu}^{S,RR}\big{]}_{1111},\big{[}{C}_{e(u)\gamma}\big{]}_{11}$
Im$(y_{11}^{e}y_{22}^{u})$ $(-2\pm 3)\times 10^{-11}$ $(-2\pm 3)\times 10^{-12}$ $\big{[}{C}_{e(u)\gamma}\big{]}_{11}$, $\big{[}{C}_{\tilde{G}}\big{]}$
Table 12: Current and future EDM constraints on the combinations of the generic flavor conserving Higgs couplings. In column 3, we also indicate the origin of leading contributions to the EDMs at $\mu=2$GeV. Its worth mentioning that the EDMs can also probe other combinations of the Yukawa couplings such as $y_{ii}^{f}y_{jj}^{f^{\prime}*}$, for $f=f^{\prime}$ or $f\neq f^{\prime}$, which are not shown here. In that case, the scalar or vector operators with left-right chiralities are generated at the EW scale (see e.g. (127)). The constraints on the leptonic and semileptonic operators are dominated by the electron EDM.
@@ 6.4 LHC constraints on flavor-conserving Higgs couplings
where $\sigma^{\rm SM}$ and $\sigma$ are, respectively, the production cross sections in the SM and the correction induced by non-standard interactions. $\Gamma^{\rm SM}_{H\rightarrow f}$ are the decay widths in the channel $f$ and $\Gamma^{\rm SM}_{\textrm{tot}}$ the Higgs total width. In the SM, the Yukawas of the light $u$, $d$, and $s$ quarks are too small to significantly affect Higgs production and decay. Direct constraints on $H\rightarrow c\bar{c}$ are sensitive to charm Yukawas about ten times larger than the SM expectation ATLAS:2022ers ; CMS:2022fxs , while the $b$ and $t$ Yukawas are compatible with the SM ATLAS:2022vkf ; CMS:2022dwd . In the HEFT, the quark masses and Yukawas are decoupled from each other. Light quark Yukawas can thus affect the Higgs production via the partonic processes $\bar{q}q\rightarrow H$, which falls in the same category at gluon fusion, and $q\bar{q}^{(\prime)}\rightarrow HV$, with $V=W,Z$. They in addition affect the decay signal strength by modifying the Higgs total width. The top Yukawa can be probed by the loop processes $gg\rightarrow H$, $H\rightarrow\gamma\gamma$, $H\rightarrow\gamma Z$, and, at tree level, via $t\bar{t}H$ production.
@@
The Yukawa also affect the Higgs decay into fermions. At NLO in QCD Spira:2016ztx
@@
For illustration, we use the signal strength measurements of the ATLAS experiment ATLAS:2022vkf , set $\kappa_{V}=1$ and $\textrm{Re}\,y_{t}$ to its SM value. The latter assumption is justified by analyses of the CP properties of the couplings of the Higgs boson to the top, which prefer a dominantly real coupling CMS:2020cga ; ATLAS:2020ior . With these assumptions, a combined fit to the Yukawas of the electron and of the $u$, $d$, $s$, $c$ and $b$ quarks to ATLAS data yields
$|y^{e}_{11}|$ $<1.3\cdot 10^{-2},\qquad|y^{u}_{11}|<6.6\cdot 10^{-3},\qquad|y^{d}_{11}|<7.0\cdot 10^{-3},$ (142)
$\quad|y^{d}_{22}|$ $<7.3\cdot 10^{-3},\qquad|y^{u}_{22}|<7.3\cdot 10^{-3},$ (143)
@@
We next perform a joint fit to the Higgs and EDM data in the HEFT framework, in which fermion masses and Yukawa are independent. In our fit, we fix $\kappa_{V}$ and the real part of the top Yukawa coupling, $\textrm{Re}y^{u}_{33}$, to their SM model values. We then perform a simultaneous fit to the real parts of the $e$, $u$, $d$, $s$, $c$ and $b$ Yukawas and to one imaginary part at a time. At the 90% CL, we find
$|\textrm{Im}y^{u}_{11}|$ $<1.4\cdot 10^{-6},\qquad|\textrm{Im}y^{d}_{11}|<6.0\cdot 10^{-7},\qquad|\textrm{Im}y^{d}_{22}|<7.4\cdot 10^{-3},$ (145)
$|\textrm{Im}y^{u}_{22}|$ $<1.65\cdot 10^{-3},\qquad|\textrm{Im}y^{d}_{33}|<1.2\cdot 10^{-2}.$ (146)
We see that for the $u$, $d$ and $c$ Yukawas, the bound on the imaginary part is much stronger than the Higgs bounds in Eq. (142). For the $s$ and $b$ quark, however, the bounds go back to what we obtained from the signal strengths.

### v2

@@ 6.2 Constraints on flavor-violating Higgs couplings
$\hfil\displaystyle d_{\rm Hg}\over[{\rm ecm}]$ $\supset{(1.1\pm 4.1)\cdot 10^{-24}\over(2\times 0.125^{2})}{\rm Im}(y^{d}_{12}y^{d}_{21})+{(-5.7\pm 34)\cdot 10^{-24}\over(2\times 0.125^{2})}{\rm Im}(y^{d}_{13}y^{d}_{31})$ (133)
$+{(-6.4\pm 4.7)\cdot 10^{-28}\over(2\times 0.125^{2})}{\rm Im}(y^{d}_{23}y^{d}_{32})+{(-2.4\pm 15.3)\cdot 10^{-24}\over(2\times 0.125^{2})}{\rm Im}(y^{u}_{12}y^{u}_{21})\,,$
@@
${\omega_{\textrm{HfF}}\over[\textrm{mrad/s}]}\supset{(9.1\pm 0.4)\cdot 10^{5}\over(2\times 0.125^{2})}{\rm Im}(y^{e}_{12}y^{e}_{21})+{(1.5\pm 0.1)\cdot 10^{7}\over(2\times 0.125^{2})}{\rm Im}(y^{e}_{13}y^{e}_{31}).$ (134)
Similarly, one can find expressions for the other EDM observables. Only a specific combination of the Higgs coupling can be probed with EDMs. To get an estimate of the upper bounds on these couplings one can compare the above expressions with the experimental limits, e.g. for $\left|d_{\rm Hg}\right|<6.2\times 10^{-30}$ $e$ cm and $|\omega_{\textrm{HfF}}|<0.17$ mrad/s.
Figure 5: EDM and flavor constraints on the CP and flavor violating Higgs couplings. Note that the black line, brown and golden lines are truncated to increase the statistics.
We also perform a combined fit of FV Higgs couplings to all EDM observables ($d_{n}$, $d_{\rm Hg}$, $d_{\rm Xe}$, $d_{\rm Ra}$, $\omega_{\rm HfF},\omega_{\rm YbF},\omega_{\rm ThO}$) using the current and future upper limits given in Tab. 2. For this purpose, we assume only one combination of the couplings to be non-zero at a time. The results are presented in Tab. 11 (column 2).
In column 3, we indicate the origin of the EDMs in terms of low-energy WET operators at $\mu=2$ GeV. At the EW scale, the scalar and vector operators are generated by integrating out the Higgs as shown in the matching relations (127). Some of these give a leading contribution to the EDMs through direct four-fermion matrix elements (see (3.4.1)- (60)) for the light flavors $f,f^{\prime}=u,d,s$. 1-loop QCD and QED effects through the operator mixing onto dipoles are also present for the light flavors. However, for the heavier generations, $c$ and $b$, the contribution to the EDMs arises solely due to operator mixing at 1-loop level (see (80)-(82)). Apart from that, as shown in column 3, the Weinberg operator also contributes for the case of heavier generations due to threshold corrections of the $b$ and $c$ dipoles onto it at the $m_{b}$ and $m_{c}$ scales (see Eqs. (89) and (90)).
@@
Im($y^{d}_{12}y_{21}^{d}$) $1.2\cdot 10^{-7}$ $7.0\cdot 10^{-9}$ $\big{[}{C}_{d\gamma(G)}\big{]}_{11,22},\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{1221},\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{1122}$
Im($y_{13}^{d}y_{31}^{d}$) $2.1\cdot 10^{-8}$ $6.0\cdot 10^{-10}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{d\gamma(G)}\big{]}_{11}$
Im($y_{23}^{d}y_{32}^{d}$) $1.5\cdot 10^{-4}$ $1.4\cdot 10^{-6}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{d\gamma(G)}\big{]}_{22}$
Im($y_{12}^{u}y_{21}^{u}$) $5.0\cdot 10^{-8}$ $2.3\cdot 10^{-9}$ $\big{[}{C}_{\widetilde{G}}\big{]},\big{[}{C}_{u\gamma(G)}\big{]}_{11}$
Im($y_{12}^{e}y_{21}^{e}$) $1.1\cdot 10^{-9}$ $1.1\cdot 10^{-10}$ $\big{[}{C}_{e\gamma}\big{]}_{11}$
Im($y_{13}^{e}y_{31}^{e}$) $6.0\cdot 10^{-11}$ $6.0\cdot 10^{-12}$ $\big{[}{C}_{e\gamma}\big{]}_{11}$
Table 11: Current and future EDM constraints on the combinations of the generic flavor violating Higgs couplings (column 2) at $1\sigma$ level . In the column 3, we also indicate the origin of EDM contributions at $\mu=2$ GeV. The matching scale is set to the Higgs mass.
@@ 6.3 EDM constraints on flavor-conserving Higgs couplings
For the flavor-conserving Yukawas, we get contributions from the tree-level matching onto four-fermion, given in Eq. (127), and from the 2-loop Barr-Zee diagrams, in Eq. (6.1). Using these one can directly use the master formulae to obtain the EDMs in terms of the FC Yukawas, similar to the expressions in (133) and (134).
@@
here $f$ can be equal to $f^{\prime}$. Therefore, for $f=f^{\prime}$ and $i=j$, one can also probe individual couplings (more precisely a combination of its Re and Im parts: $\textrm{Re}(y^{f}_{ii})\textrm{Im}(y^{f}_{ii})$) in an uncorrelated manner to other couplings, under the assumption that only a single coupling is turned on. Overall, the tree-level and 2-loop BZ diagram effects can test very different parts of the low-energy Higgs effective coupling parameter space. In Tab. 12 we show upper bounds at $1\sigma$ level on the various combinations of the generic FC Higgs couplings.
@@
Im$(y_{11}^{d}y_{11}^{d})$ $2.4\cdot 10^{-8}$ $2.6\cdot 10^{-9}$ $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{1111},\big{[}{C}_{d\gamma(G)}\big{]}_{11}$
Im$(y_{22}^{d}y_{22}^{d})$ $2.0\cdot 10^{-3}$ $1.9\cdot 10^{-5}$ $\big{[}{C}_{dd}^{S1(8),RR}\big{]}_{2222},\big{[}{C}_{d\gamma(G)}\big{]}_{22}$
Im$(y_{33}^{d}y_{33}^{d})$ $1.1\cdot 10^{-3}$ $9.0\cdot 10^{-6}$ $\big{[}{C}_{\widetilde{G}}\big{]}$
Im$(y_{11}^{u}y_{11}^{u})$ $2.6\cdot 10^{-8}$ $1.6\cdot 10^{-9}$ $\big{[}{C}_{uu}^{S1(8),RR}\big{]}_{1111},\big{[}{C}_{u\gamma(G)}\big{]}_{11}$
Im$(y_{22}^{u}y_{22}^{u})$ $4.0\cdot 10^{-4}$ $3.5\cdot 10^{-6}$ $\big{[}{C}_{\widetilde{G}}\big{]}$
Im$(y_{11}^{e}y_{11}^{e})$ $1.2\cdot 10^{-7}$ $1.2\cdot 10^{-8}$ $\big{[}{C}_{ee}^{S,RR}\big{]}_{1111},\big{[}{C}_{e\gamma}\big{]}_{11}$
Table 12: Current and future EDM constraints on the combinations of the generic flavor conserving Higgs couplings at $1\sigma$ level (to obtain $90\%$CL these numbers have to be multiplied by factor 1.6 ). In column 3, we also indicate the origin of leading contributions to the EDMs at $\mu=2$GeV. It is worth mentioning that the EDMs can also probe other combinations of the Yukawa couplings which are not shown here. In that case, the scalar or vector operators with left-right chiralities are generated at the EW scale (see e.g. (127)). The electron EDM dominates the constraints on the leptonic and semileptonic operators. The matching scale is set to the Higgs mass.
@@ 6.4 LHC constraints on flavor-conserving Higgs couplings
where $\sigma^{\rm SM}$ and $\sigma$ are, respectively, the production cross sections in the SM and the correction induced by non-standard interactions. $\Gamma^{\rm SM}_{H\rightarrow f}$ are the decay widths in the channel $f$ and $\Gamma^{\rm SM}_{\textrm{tot}}$ the Higgs total width. In the SM, the Yukawas of the light $u$, $d$, and $s$ quarks are too small to significantly affect Higgs production and decay. Direct constraints on $H\rightarrow c\bar{c}$ are sensitive to charm Yukawas about ten times larger than the SM expectation ATLAS:2022ers ; CMS:2022fxs , while the $b$ and $t$ Yukawas are compatible with the SM ATLAS:2022vkf ; CMS:2022dwd . In the HEFT, the quark masses and Yukawas are decoupled from each other. Light quark Yukawas can thus affect the Higgs production via the partonic processes $\bar{q}q\rightarrow H$, which falls in the same category at gluon fusion, and $q\bar{q}^{(\prime)}\rightarrow HV$, with $V=W,Z$. They also affect the decay signal strength by modifying the Higgs total width. The top Yukawa can be probed by the loop processes $gg\rightarrow H$, $H\rightarrow\gamma\gamma$, $H\rightarrow\gamma Z$, and, at tree level, via $t\bar{t}H$ production.
@@
The Yukawa also affect the Higgs decay into fermions. At NLO in QCD, the decay of a Higgs boson to two quarks is given by Spira:2016ztx
@@
The decay into two leptons has a similar expression, without the $\mathcal{O}(\alpha_{s})$ correction. For illustration, we use the signal strength measurements of the ATLAS experiment ATLAS:2022vkf , set $\kappa_{V}=1$ and $\textrm{Re}\,y^{u}_{33}$ to its SM value. The latter assumption is justified by analyses of the CP properties of the couplings of the Higgs boson to the top, which prefer a dominantly real coupling CMS:2020cga ; ATLAS:2020ior . With these assumptions, a combined fit to the Yukawas of the electron and of the $u$, $d$, $s$, $c$ and $b$ quarks to ATLAS data yields
$|y^{e}_{11}|$ $<1.2\cdot 10^{-2},\qquad|y^{u}_{11}|<6.4\cdot 10^{-3},\qquad|y^{d}_{11}|<6.9\cdot 10^{-3},$
$\quad|y^{d}_{22}|$ $<7.1\cdot 10^{-3},\qquad|y^{u}_{22}|<7.1\cdot 10^{-3},$ (142)
@@
We next perform a joint fit to the Higgs and EDM data in the HEFT framework, in which fermion masses and Yukawa are independent. In our fit, we fix $\kappa_{V}$ and the real part of the top Yukawa coupling, $\textrm{Re}y^{u}_{33}$, to their SM model values. We then perform a simultaneous fit to the real parts of the $e$, $u$, $d$, $s$, $c$ and $b$ Yukawas and to one imaginary part at a time. We neglect theoretical errors on EDMs, so that the bounds give an indication of the full potential of EDMs experiments, but should be taken with some caution. If we include both the Barr-Zee and four-fermion contributions, at the 90% CL we find
$|\textrm{Im}y^{e}_{11}|<1.2\cdot 10^{-6},\qquad|\textrm{Im}y^{u}_{11}|$ $<1.2\cdot 10^{-6},\qquad|\textrm{Im}y^{d}_{11}|<1.7\cdot 10^{-7},$
$|\textrm{Im}y^{d}_{22}|<7.1\cdot 10^{-3},\qquad|\textrm{Im}y^{u}_{22}|$ $<7.1\cdot 10^{-3},\qquad|\textrm{Im}y^{d}_{33}|<1.2\cdot 10^{-2}.$ (144)
The 90% CL for the real part of the couplings corresponds to the ranges given in Eqs. (6.4) and (143). We see that for the $e$, $u$ and $d$ Yukawas, the EDM bound on the imaginary part is much stronger than the bounds from Higgs observables in Eq. (6.4). For the $s$, $c$ and $b$ quark, however, the bounds go back to what we obtained from the signal strengths. This is partially due to cancellations between the Barr-Zee (Eq. (6.1)) and four-fermion (Eq. (127)) contributions to EDMs, which can happen once we let the real part of the Yukawas deviate from its SM value.